OpenStudy (anonymous):

What's 2+2 funniest answer gets a medal

OpenStudy (anonymous):

4 i think

OpenStudy (anonymous):

funnier i dont think

OpenStudy (anonymous):

That's funny, batman standards

OpenStudy (anonymous):

Also, 2+2=11

OpenStudy (anonymous):

Fish durrr.... this is considered spam ~ plz dont post question like this

OpenStudy (the_fizicx99):

Baman pls, it's 22.

OpenStudy (anonymous):

how is it 11? and its not a spam

OpenStudy (anonymous):

http://i.imgur.com/hinAw.gif Me waiting, till a mod deletes this.

OpenStudy (anonymous):

2+1 = 10, 2 + 2 = 2+1+1 = 10+1 = 11

OpenStudy (anonymous):

okay wio wins

OpenStudy (anonymous):

@iambatman Where is the signal?

OpenStudy (the_fizicx99):

Lies ;-;

OpenStudy (anonymous):

/.-' yes it is considered spam ~ CoC

OpenStudy (anonymous):

and u know guys try having some fun next time don't be so nerdy

OpenStudy (the_fizicx99):

Wio your answer would actually be 20 >.<

OpenStudy (anonymous):

come on get out now

OpenStudy (the_fizicx99):

Using your imagination there's some sense to what wio did cx keyword imagination. Don't judge based on my thought >.>